OpenStudy (anonymous):

A motorboat travels 273km in 7 hours going upstream. It travels 399km going downstream in the same amount of time. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (anonymous):

please give me the answer im one away

OpenStudy (dtan5457):

You got the last question you asked?

OpenStudy (dtan5457):

Ok, well to start this you need the rates of both streams

OpenStudy (dtan5457):

for the upstream one you do 273/7=39km an hour downstream: 399/7=57km an hour downstream is the one with the current so you need now a system let me represent the boat and c represent the current b+c=57km b-c=39km add them to solve for b 2b=96 b=48km<<<speed of the boat 48+c=57km c=9km<<speed of current

OpenStudy (dtan5457):

by still water you will refer to the speed of the boat aka 57km

OpenStudy (perl):

Let x = speed of boat in still water y = speed of water current Use equation distance = rate * time 273 = (x - y)*7 399 = ( x + y) * 7

OpenStudy (anonymous):

i got that but it said i was wrong

OpenStudy (dtan5457):

*48 km

OpenStudy (dtan5457):

not sure what i did wrong. it looks right to me.